Laura Dawson can get to school in 1/4 hour if she rides her bike. It takes her 3/4 if she walks. Her speed when walking is 6 miles per hour slower than her speed when riding.What is her speed when riding?
Answer by stanbon(75887)

Laura Dawson can get to school in 1/4 hour if she rides her bike.
It takes her 3/4 hr. if she walks.
Her speed when walking is 6 miles per hour slower than her speed when riding.
----
Walking DATA:
time = 3/4 hr ; speed = r mph ; distance = (3/4)r miles
--------
Riding DATA:
time = 1/4 hr ; speed = r+6 mph ; distance = (1/4)(r+6)
------
Equation:
distance = distance
(3/4)r = (1/4)(r+6)
3r = r+6
2r = 6
r = 3 mph (walking speed)
What is her speed when riding?
r+6 = 9 mph (riding speed)
==============================
Cheers,
Stan H.
=============